### v1

@@ 4.3 Ground-Truth Labeling
Because our mechanistic approach requires analyzing the model’s unique generation trace, we must label the model’s own output. Given this constraint and the need for a scalable solution, we implemented an automated LLM-as-a-judge approach. Recognizing that the validity of LLM-based labels is crucial, we implemented a rigorous, multi-stage validation process for our LLM-as-a-judge (Llama-3.1-70B-Instruct):
  • Constrained task: We tasked the judge with a narrow attestation judgment using the ARGUE framework [32], not an open-ended quality evaluation. This limits subjectivity and improves reliability.
  • Stability testing: We verified the reliability of the judgments by testing for stability across multiple runs at varying temperatures. The labels were highly consistent, with less than 1.5% of citations showing any variation.
  • Human validation study: To verify the accuracy of the LLM-as-a-judge, we conducted a formal validation study using Cohen’s Kappa ($\kappa$) on a subset of 100 randomly sampled judgments (50 from Llama-3.1-8B and 50 from Llama-3.2-3B) using two independent human assessors [33]. As shown in Table 1, the human annotators achieved substantial agreement ($\kappa=0.68$). Crucially, the LLM judge demonstrated high alignment with both human raters across both model scales, confirming its reliability as a proxy for human judgment in this study.
Table 1: Validation study results using Cohen’s Kappa ($\kappa$). Scores indicate inter-rater reliability (Human-Human) and LLM-as-a-judge alignment against human ground truth.
Model Inter-Rater Human-LLM Alignment Overall
(Target) (H1 vs. H2) (LLM vs. H1) (LLM vs. H2) LLM Mean
Llama-3.1-8B 0.79 0.67 0.61 0.64
Llama-3.2-3B 0.62 0.62 0.79 0.71
Average 0.71 0.65 0.70 0.68
@@ 4.4 Models and Data Statistics
We test our framework on two state-of-the-art long-context models: Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct. The 8B model generated longer reports on average ($\approx$ 1,091 tokens) than the 3B model ($\approx$ 782 tokens). For the 3B model, our final dataset contains 208 correct and 93 hallucinated citation tokens. For the 8B model, we identified 465 correct and 107 hallucinated citation tokens. To train our classifiers, we use a balanced training set and evaluate on a held-out, imbalanced test set that mirrors the real-world distribution of correct vs. hallucinated citations.
@@ 4.5 Feature Engineering and Classification
The raw mechanistic scores from our framework produce a high-dimensional feature vector for each citation token. For the 32-layer, 32-head Llama-3.1-8B model, a single score such as CAS generates 1,024 distinct measurements (one for each head at each layer) for every citation token we analyze. The high dimensionality of these raw features relative to the number of labeled samples creates a significant risk of overfitting. To mitigate this, we developed a multi-stage pipeline to distill these raw signals into a concise and robust feature set.
First, we implement model component pruning. We hypothesize that many layers and heads are not specialized for citation attribution. Therefore, we rank all model components (i.e., each head at each layer) for each mechanistic score based on their individual correlation with the citation hallucination label. We treat the pruning percentage, $k$, as a hyperparameter to select only the top-$k\%$ model components from each of our mechanistic scores independently. This approach is agnostic to a model component’s architectural location, allowing us to discover whether the most predictive signals are distributed broadly or are localized to a specific subset of model components. Our optimization revealed that the optimal $k$ varies by model size; we report final results using the optimal $k$ for each model ($k=100\%$ for 3B, $k=25\%$ for 8B).
Second, the pruned features undergo dimensionality reduction. For attention-based scores (CAS, BAS, and the baseline ECS), we aggregate across the head dimension at each layer using simple statistics (e.g., mean, standard deviation), as heads within a layer function as an unordered set. This results in a per-layer feature vector for each score. To capture trends across the model’s hierarchical layers, we engineer features using both simple aggregates (mean, std) and sequence-aware functions (linear slope, FFT) common in time-series analysis. From this candidate pool, we select the most predictive feature for each mechanistic score using in-distribution training data. This pipeline, applied identically to our FACTUM scores and the ReDeEP baselines, yields a concise final feature set; for example, our complete FACTUM model uses just four features.
These features are used to train a suite of interpretable classifiers, including Logistic Regression [49], LightGBM [18], and an Explainable Boosting Machine (EBM) [27]. We employ 10-fold stratified cross-validation for evaluation. To prevent data leakage and ensure generalizability, all data splits are performed at the report level, guaranteeing that all citations from a given report belong exclusively to either the training or the testing set within any given fold.
@@ 4.6 Baselines
To evaluate FACTUM’s effectiveness, we benchmark its performance against state-of-the-art baselines that also rely exclusively on internal model signals [^3] Code available at https://github.com/Maximeswd/citation-hallucination. Our primary comparison is a re-implementation of the mechanistic scores from ReDeEP: the External Context Score (ECS) and the Parametric Knowledge Score (PKS) [47]. This allows for a direct, head-to-head comparison of feature effectiveness. We also compare against a range of established confidence-based scores that probe model internals for uncertainty signals. This category includes token-level Perplexity [14], LN-Entropy [30], the Energy score [26], and the P(True) score [17].
@@ 5.1 FACTUM Establishes a New State-of-the-Art in Citation Hallucination Detection
As shown in Table 2, our results demonstrate that FACTUM significantly improves feature-based citation hallucination detection. While some classifier-free baselines achieve competitive AUC scores or high recall, this performance comes at a cost to precision, making them unreliable in practice. Consider the P(True) score for instance, which achieves high recall on the 8B model. While it successfully identifies a large proportion of hallucinated citations (low false negatives), it does so at the cost of fairly poor precision. This means it incorrectly flags a vast number of correct citations as hallucinations (a high false positive rate). For any practical application, such a system would be unusable, as it would constantly undermine user trust. This highlights the critical need for a more balanced and nuanced classification approach that can reliably distinguish between correct and hallucinated citations without overwhelming the user with false alarms.
[…]
Table 2: Mean performance on the NeuCLIR 2024 dataset. The table is split into two sections: classifier-free General Baselines and our main classifier-based results. In the top section, the best general baseline for each metric is marked with a gray background. In the bottom section, highlighting is scoped within each classifier block, with the best (bold) and runner-up (underlined) performance shown. A star () indicates a statistically significant improvement over ‘ECS+PKS (Baseline)’ ($p_{FDR}<0.05$). A dagger () indicates that the improvement of ‘FACTUM (Ours)’ over ‘CAS+V_ffn (Ours)’ is statistically significant.
@@ 5.2 Uncovering the Mechanistic Signatures of Citation Hallucination
Beyond detection, FACTUM serves as a diagnostic tool to uncover the internal signatures of citation. We initially hypothesized that correct citations would be characterized by higher contextual alignment, more utilization of the beginning-of-sentence token as an attention sink, higher parametric force, and higher pathway alignment. The results in Table 3 demonstrate that the 3B model using 100% of its model components strongly supports this: Correct citations are indeed marked by significantly higher CAS, BAS, PAS, and PFS. This indicates that at this scale, utilizing all model components, the most effective strategy for correct citation generation is one of direct reinforcement, where all internal mechanisms work in concert.
In contrast, the 8B model’s signature reveals a shift toward mechanistic specialization. Its peak performance requires pruning to its top 25% of model components, suggesting that the task is localized to a more specialized set of internal mechanisms. Within this subset, while the constructive roles of high PFS and BAS remain consistent, CAS becomes inconsistent, and PAS becomes significantly lower for our best-performing classifier (Logistic Regression). This suggests that within this specialized group of model components, the pathways are not required to simply agree, but to contribute distinct, orthogonal information to the residual stream.
This forces a refinement of our hypothesis: as models scale, and don’t use all their model components, their strategy for correct citation appears to evolve from simple “agreement” to “specialized contribution”, where pathways provide complementary, orthogonal information. FACTUM’s ability to reveal these emergent, scale-dependent strategies proves that a “one-size-fits-all” theory of citation hallucination is insufficient.
Table 3: Mechanistic signatures of correct vs. hallucinated citations. We performed a one-sided Mann-Whitney U test on the features from the ‘Full FACTUM (Ours)’ set. Cells are colored to show the direction of the significant relationship: $\uparrow$ indicates the score was significantly higher for correct citations, while $\downarrow$ indicates it was significantly lower ($p<0.05$).
@@
BAS (Attention Sink Usage) $\uparrow***$ $\uparrow*$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$
CAS (Contextual Alignment) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\downarrow***$
PAS (Pathway Alignment) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow**$ $\uparrow**$ $\downarrow***$
PFS (Parametric Force) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$
*$p<0.05$, **$p<0.01$, ***$p<0.001$.
@@ 5.3 Case Study: Mechanistic Analysis of Figure 1
To illustrate the diagnostic utility of our framework, we analyze the transition from correct to hallucinated citation in Figure 1 through the interplay of our four mechanistic signals.
1) Characterizing Attributional Drift. In the correct scenario (Left), the internal states reflect a high degree of mechanistic correspondence between the claim and its origin in Document 1. In the hallucinated scenario (Right), the model’s internal signature is consistent with a vulnerability to attributional drift. Mechanistically, this can be viewed as a divergence between information streams: while the FFN pathway continues to apply significant parametric force (PFS) towards the correct factual claim (100 mg), the Attention pathway fails to ground the citation digit in the source documents (Low CAS). Furthermore, a decrease in BAS suggests a failure in the internal synthesis mechanism required to manage information flow over longer contexts. This shift indicates that the generated citation index is likely driven by parametric likelihood or local sequence patterns rather than verified contextual grounding.
2) Coordination vs. Dissonance. FACTUM identifies these failures by capturing the holistic relationship between pathways at the moment of citation:
  • Pathway Coordination (Correct Case): When the model cites Document 1, FACTUM identifies coordination across all signals. The Attention pathway retrieves the claim from the correct context (High CAS) and utilizes the attention sink for stable synthesis (High BAS). Simultaneously, the FFN pathway provides a corroborating update (High PFS) that is semantically coordinated with the retrieved evidence (High PAS). Here, the reader and recall mechanisms work in concert to support the generated token.
  • Mechanistic Dissonance (Hallucinated Case): When the model falsely cites Document 2, a systemic dissonance emerges. Because Document 2 contains conflicting information (10 mg), the Attention pathway exhibits low semantic alignment with that source (Low CAS) and reduced synthesis stability (Low BAS). Crucially, the FFN pathway continues to prioritize the “100 mg” fact (Sustained PFS), resulting in an update that is orthogonal or contradictory to the cited context (Low/Divergent PAS).
By identifying this internal dissonance, FACTUM serves as a potential verifiability guardrail. As Table 3 demonstrates, correct citations are rarely defined by a single metric; rather, it is the systemic coordination of grounding, synthesis, force, and alignment that differentiates verifiable attribution from mechanistic drift, a signature entirely invisible to black-box evaluation methods.
6 Conclusion
We introduced FACTUM, a framework that establishes a new state-of-the-art method for detecting citation hallucination by analyzing a model’s internal pathways. As detailed above, its key contribution is revealing the evolving, scale-dependent signatures of correct attribution. These findings challenge prevailing assumptions about the role of parametric knowledge and demonstrate the need for nuanced, mechanistic tools to build more reliable and trustworthy RAG systems.
A Final Note on Trust of Citations.
@@ 6 Conclusion
A question for the reader: Did you notice our citation of Shakespeare in the introduction to support a claim about how citations incite trust? Although the Sonnets explore themes of trust, making the reference seem plausible, it was a deliberate false citation. Now imagine if the citation appeared as: “…incite trust in the reader [43]*”. That asterisk would represent a potential fix: a visual warning generated by a system like FACTUM, indicating the citation’s grounding is highly suspect.
We included this small deception to make the central point of this paper tangible: even for experts reading a paper on this exact topic, the pull to trust a citation is powerful. This subtle gap between plausibility and verifiable truth is precisely highlighting the critical need for our robust, mechanistic detection method FACTUM.
Acknowledgments

### v2

@@ 4.3 Ground-Truth Labeling
Because our mechanistic approach requires analyzing the model’s unique generation trace, we must label the model’s own output. Given this constraint and the need for a scalable solution, we implemented an automated LLM-as-a-judge approach using Llama-3.1-70B-Instruct. Recognizing that the validity of LLM-based labels is crucial, we implemented a rigorous, multi-stage validation process. First, we constrained the judge to narrow attestation judgments using the ARGUE framework [32] rather than open-ended evaluation. Second, stability testing across multiple temperatures showed highly consistent labels ($1.5\%$ variation). Finally, we conducted a human validation study on 100 randomly sampled judgments (50 per model) using two independent assessors. Assessors achieved substantial agreement (Avg. $\kappa=0.71$), and the LLM judge aligned highly with human ground truth, achieving a mean $\kappa=0.68$ (0.64 for 8B and 0.71 for 3B). These results confirm the LLM judge’s reliability as a proxy for human judgment, providing a robust foundation for our analysis.
@@ 4.4 Models and Data Statistics
We select two models with 128k context windows to evaluate FACTUM in long-form settings: Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct. The 8B model generated longer reports on average ($\approx$ 1,091 tokens) than the 3B model ($\approx$ 782 tokens). For the 3B model, our final dataset contains 208 correct and 93 hallucinated citation tokens. For the 8B model, we identified 465 correct and 107 hallucinated citation tokens. To train our classifiers, we use a balanced training set and evaluate on a held-out, imbalanced test set that mirrors the real-world distribution of correct vs. hallucinated citations.
@@ 4.5 Feature Engineering and Classification
The raw mechanistic scores from our framework produce high-dimensional feature vectors per citation token. For the 32-layer, 32-head Llama-3.1-8B model, a single score such as CAS generates 1,024 distinct measurements (one for each head at each layer) for every citation token we analyze. To mitigate overfitting risks from such high dimensionality relative to sample size, we developed a multi-stage pipeline to distill these signals into a concise feature set.
First, we prune model components, because we hypothesize that citation attribution may be localized to specialized layers and heads. Therefore, we rank all model components (i.e., each head at each layer) for each mechanistic score based on their individual correlation with the citation hallucination label using only the training set to prevent leakage. We treat the pruning percentage, $k$, as a hyperparameter to select the top-$k\%$ components independently for each mechanistic score. This approach is agnostic to architectural location, allowing us to discover whether predictive signals are distributed broadly or localized to a specialized subset. Our optimization revealed that the optimal $k$ varies by model size, suggesting significant specialization at scale; we report final results using the optimal $k$ for each model ($k=100\%$ for 3B, $k=25\%$ for 8B).
Second, the pruned features undergo dimensionality reduction. For attention-based scores, we aggregate across the head dimension at each layer (e.g., mean, standard deviation), as heads within a layer function as an unordered set. This results in a per-layer feature vector for each score. To capture trends across the model’s hierarchical layers, we engineer features utilizing sequence-aware functions (linear slope, FFT) common in time-series analysis [11]. From this candidate pool, we select the most predictive feature for each mechanistic score using in-distribution training data. This pipeline, applied identically to our FACTUM scores and the ReDeEP baselines, yields a concise final feature set; for example, our complete FACTUM model uses just four features.
Finally, we train interpretable classifiers, including Logistic Regression [48], LightGBM [19], and an Explainable Boosting Machine (EBM) [27] using 10-fold stratified cross-validation. To prevent data leakage, all splits are performed at the report level, ensuring all citations from a given report belong exclusively to either the training or the testing set within any fold.
@@ 4.6 Baselines
To evaluate FACTUM, we benchmark its performance against state-of-the-art baselines that also rely exclusively on internal model signals [^3] Code available at https://github.com/Maximeswd/citation-hallucination. Our primary comparison is a re-implementation of the mechanistic scores from ReDeEP: the External Context Score (ECS) and the Parametric Knowledge Score (PKS) [46]. This allows for a direct, head-to-head comparison of feature effectiveness. We also compare against a range of established confidence-based scores that probe model internals for uncertainty signals. This category includes token-level Perplexity [15], LN-Entropy [30], the Energy score [26], and the P(True) score [18].
@@ 5.1 FACTUM Establishes a New State-of-the-Art in Citation Hallucination Detection
As shown in Table 1, our results demonstrate that FACTUM significantly improves feature-based citation hallucination detection. While some classifier-free baselines achieve competitive AUC or recall, poor precision makes them unreliable in practice. For example, P(True) achieves high recall on the 8B model, but incorrectly flags numerous correct citations as hallucinations (a high false positive rate), undermining user trust. This highlights the need for a balanced and nuanced classification approach that can reliably distinguish between correct and hallucinated citations without overwhelming the user with false alarms.
Among the classifier-based methods, our analysis reveals a clear hierarchy: FACTUM (Ours) >> CAS + PFS (Ours) >> ECS + PKS (Baseline). Firstly, our refined CAS + PFS (Ours) significantly outperforms the ReDeEP baseline. This confirms the superior signal quality of direct pathway scores that avoid the noise of LogitLens and prompt-level attention dilution. Second, the full FACTUM framework achieves the highest overall performance, winning 27 of 30 head-to-head comparisons with a peak AUC of 0.737 with the Logistic Regression classifier. Crucially, FACTUM increases 8B model precision from 0.201 to 0.334, which is a 66% relative improvement over the ReDeEP baseline. This statistically significant improvement of the full FACTUM framework over the already strong “CAS + PFS (Ours)” (indicated by the $\dagger$} symbol), validates our central hypothesis: that a complete and reliable detection framework must account not only for the primary information pathways but also for their interaction and synthesis, signals uniquely captured by our PAS and BAS scores.
Table 1: Mean performance on the NeuCLIR 2024 dataset, averaged over 10-fold stratified cross-validation. The table is split into two sections: classifier-free General Baselines and our main classifier-based results. In the bottom section, highlighting is scoped within each classifier block, with the best (bold) and runner-up (underlined) performance shown. A star () indicates a statistically significant improvement over ‘ECS+PKS (Baseline)’ ($p_{\text{FDR}}<0.05$) using a two-tailed t-test with Benjamini-Hochberg (BH) correction. A dagger ($\dagger$}) indicates that the improvement of ‘FACTUM (Ours)’ over ‘CAS+PFS (Ours)’ is statistically significant.
@@ 5.2 Uncovering the Mechanistic Signatures of Citation Hallucination
Beyond detection, FACTUM serves as a diagnostic tool to uncover the internal signatures of citation hallucination. We initially hypothesized that correct citations would be characterized by higher contextual alignment, greater attention sink usage, higher parametric force, and higher pathway alignment. Table 2 demonstrates that the 3B model using 100% of its model components strongly supports this: Correct citations are indeed marked by significantly higher CAS, BAS, PAS, and PFS. This indicates that at this scale, utilizing all model components, the most effective strategy for correct citation generation is one of direct reinforcement, where all internal mechanisms work in concert.
In contrast, the 8B model’s signature reveals a shift toward mechanistic specialization. Its peak performance requires pruning to its top 25% of model components, suggesting that the task is localized to a specialized set of internal mechanisms. Within this subset, while the constructive roles of high PFS and BAS remain consistent, CAS becomes inconsistent, and PAS becomes significantly lower for our best-performing classifier (Logistic Regression). This suggests that within this specialized group of model components, the pathways are not required to simply agree, but to contribute distinct, orthogonal information to the residual stream. This finding relates to post-rationalized citations [50]. While one might assume models only correctly cite what they already ‘know’ (High PFS), our results show that correct citations require simultaneous grounding (CAS) and stable synthesis (BAS). In reliable RAG, parametric memory does not replace context, but actively corroborates it through orthogonal coordination.
This forces a refinement of our hypothesis: as models scale, and don’t use all their model components, their strategy for correct citation appears to evolve from simple “agreement” to “specialized contribution”, where our most accurate detector identifies pathways providing complementary, orthogonal information. The emergence of these scale-dependent strategies indicates that a “one-size-fits-all” theory of citation hallucination is insufficient.
Table 2: Mechanistic signatures of correct vs. hallucinated citations. Significance is determined via a one-sided Mann-Whitney U test on the ‘Full FACTUM (Ours)’ set with Benjamini-Hochberg correction. Heatmap shades (Dark, Medium, Light) represent significance at $p<0.001$, $p<0.01$, and $p<0.05$, respectively. Blue cells indicate higher scores for correct citations ($\uparrow$); Orange cells indicate lower scores ($\downarrow$).
@@
[…]
@@ 5.3 Case Study: Mechanistic Analysis of Figure 1
We analyze Fig. 1’s transition through our mechanistic signals. While the correct case (Left) reflects correspondence, the hallucinated scenario (Right) reveals attributional drift: the FFN pathway applies significant force (High PFS) toward the “100 mg” claim, but the Attention pathway fails to ground the citation (Low CAS). Decreased BAS further suggests a synthesis failure, indicating the citation is driven by parametric likelihood rather than verified contextual grounding.
Correct citation requires scale-dependent functional coordination. In the 3B model, this manifests as reinforcement (High PAS); in the 8B model, it shifts toward specialized, non-redundant contributions (Low PAS). Conversely, incorrectly citing Document 2 triggers mechanistic dissonance. Despite conflicting context (10 mg), the FFN pathway prioritizes its internal “100 mg” memory (Sustained PFS). This marks a universal coordination failure where the FFN pathway operates independently, providing neither the reinforcement (3B) nor the specialization (8B) required for correctness. By identifying this dissonance, FACTUM serves as a verifiability guardrail against failures entirely invisible to black-box methods.
5.3.1 Concluding Remarks.
We introduced FACTUM, a framework setting a new state-of-the-art for citation hallucination detection through internal pathway analysis. Beyond detection, our work reveals that the signature of correctness is not static, but evolves with model scale. By demonstrating that parametric force acts as a constructive corroborator rather than a mere source of hallucination, we challenge prevailing assumptions and pave the way for more nuanced, mechanistic guardrails in reliable RAG systems.
5.3.2 A Final Note on Trust of Citations.
@@
A question for the reader: Did you notice our citation of Shakespeare in the introduction to support a claim about how citations incite trust? Although the Sonnets explore themes of trust, making the reference seem plausible, it was a deliberate false citation. Now imagine if the citation appeared as: “…incite trust in the reader [42]*”. That asterisk would represent a potential fix: a visual warning generated by a system like FACTUM, indicating the citation’s grounding is highly suspect. We included this small deception to make the central point of this paper tangible: even for experts reading a paper on this exact topic, the pull to trust a citation is powerful. This subtle gap between plausibility and verifiable truth is precisely highlighting the critical need for our robust, mechanistic detection method FACTUM.
5.3.3 Acknowledgements
@@
5.3.4 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.